OpenStudy (anonymous):

John's annual salary after a raise of 15% is $45,000. What was his salary before the raise?

hartnn (hartnn):

let salary before raise =x so, x + 15%x = 45000 a% =a/100 so, can you solve : x+15x/100 = 45000

OpenStudy (anonymous):

39, 130???

hartnn (hartnn):

yes, i get the same :)

OpenStudy (anonymous):

Tanks :)

hartnn (hartnn):

wlcoms :)